OpenStudy (anonymous):

if f(x) is 5x^3 -2x and g(x) is 3x^3 what is f(g(x))

OpenStudy (anonymous):

\[f(g(x)) = 5(g(x))^3 - 2g(x) = 5(3x^3)^3 - 2(3x^3) = 5\cdot27x^9 - 6x^3\]